OpenStudy (anonymous):

equation of the tangent to curve y=e^x at x=-1

OpenStudy (anonymous):

differentiate , remembering \[\frac{d(e^x)}{dx} = e^x\] then substitute in x = -1 into dy/dx to get the gradient of the tangent you should get m = e^(-1) = 1/e now use the formula y - y1 = m(x - x1) with x1 as -1 and y1 as e^(-1) = 1/e

OpenStudy (anonymous):

yes tht is the answer. I ws about to post